# The nature of mathematical models

### v1

@@ Abstract
Modeling has become a widespread, useful tool in mathematics applied to diverse fields, from physics to economics to biomedicine. Practitioners of modeling may use algebraic or differential equations, to the elements of which they attribute an intuitive relationship with some relevant aspect of reality they wish to represent. More sophisticated expressions may include stochasticity, either as observation error or system noise. However, a clear, unambiguous mathematical definition of what a model is and of what is the relationship between the model and the real-life phenomena it purports to represent has so far not been formulated. The present work aims to fill this gap, motivating the definition of a mathematical model as an operator on a Hilbert space of random variables, identifying the experimental realization as the map between the theoretical space of model construction and the computational space of statistical model identification, and tracing the relationship of the geometry of the model manifold in the abstract setting with the corresponding geometry of the prediction surfaces in statistical estimation.
@@ 1 Introduction
Mathematical modeling is central to defining and solving problems in the physical sciences, engineering or economics. Over the past few decades it has found progressively more widespread applications also in biomedicine and other “soft” sciences. In the following we will make use, just by way of example, of naively simple problems in biology: while this choice stems from the background of the author as a biomedical modeler, the mathematical ideas are equally applicable to any of the above application fields.
@@
The initial, modeling phase thus consists in imagining possible relationships among some quantitative features of interest of the reality under study, relationships usually expressed as algebraic or differential equations of various kinds. The initial concern of the present work is to understand exactly what the variables in these equations represent. Suppose for example we wished to estimate weight from height, and we were to write, as a very first attempt to quantify the relationship of weight as a function of height, the equation
@@
Notice that at the present stage we are not dealing with any numbers at all, we are at the first stage of idealization of a relationship between a generic concept of height with a generic concept of weight. Our basic question is then: what, in the above equation, is $H$? The immediate, naive answer would be "Height!". Yes, clearly, but whose height? Obviously $H$ is not meant to represent the numerical value of the height of any single individual, to be put in relationship with the weight $W$ of that same individual: changing individuals would then in general change the value of the parameter $k$, which is instead intended to express some general proportion between the height of people and their weight. Similarly, Eq. 1 does not express the relationship between an algebraic variable $W$ and an algebraic variable $H$: we do not imagine people’s weights to follow exactly that functional relationship. Rather, we might consider $H$ as indicating a quantity (height) which could be measured on some individual or on a group of individuals, quite possibly with variability due to heterogeneity among people and/or to measurement error: while we have not yet done so at the present stage of the modeling exercise, we could do so in the future. We are thus led to considering $H$ as a random variable.
@@
At this point we consider what $W$ might be. Notice that, if we were to consider hypothetically the observable weight of individuals, and, following the previous line of reasoning, we were to consider such observable weight of individuals as a random variable, this random variable (let us denote it provisionally by $W^{o}$) would not be the same as $W$. That $W$ is a random variable follows from the fact that it is a function of some numerical value $k$ and of another random variable $H$, so yes, $W$ is indeed a random variable. But $W$ is that random variable obtained from $H$ through some sort of mathematical manipulation, whereas $W^{o}$ is another random variable, representing some characteristic of the considered people, which may or may not depend on $H$ and even if it did it may well in general not depend on it through Eq. 1. In other words, we might consider $W^{o}$ as the objective (hence the o) random variable that we are trying to approximate through some other random variable $W$, which is itself computable from $H$. We are thus led to the definition of a model, which will be formalized in the following, as an operator yielding a random variable (in this case $W$) with which we try to approximate another, objective random variable(in this case $W^{o}$). As will be made more precise in the following, approximation has meaning within a space endowed at least with a norm, while other useful concepts like projections will need further structure (inner products), thus leading us to deal with of a Hilbert space of random variables.
@@
If we now consider the second, identification phase of modeling practice, we see that it is commonly performed through the minimization of a cost functional (e.g. by Ordinary Least Squares, OLS). Identification of course depends on having sampled some individuals from the population of interest and having measured on them both the intended predictor variable(s) and the intended criterion or target variable. This procedure, which is inherently computational, is executed within an $n$-dimensional Euclidean space, referred to as the case space.
@@
While both phases of the modeling process, model building and model identification, have been and are conducted innumerable times by modelers over centuries and all over the world, the relationship between ideal model building and concrete statistical estimation computations has been a matter the "obvious" intuition of the correspondence between the two. The aim of the present work is to make formally precise the relationship between the ideal construction of a mathematical model in the space of random variables and the statistical estimation procedure in $n$-dimensional case space. It should be noted that the consideration of a Hilbert space of random variables is not new, see for example Small [12]. However, the subsequent development has centered on the characterization of conditional expectation as the projection onto a linear subspace [4]. The present work, instead, focuses on the modeling process itself, giving rise to a generally nonlinear model manifold, and on the relationship of this construction with the (geometrical interpretation of) commonly performed statistical parameter estimation procedures. There is however a closer link with the whole field of information geometry [9, 11, 6], dealing with manifolds of probability distributions, in that there exists a correspondence between (equivalence classes of) random variables and their induced probability distributions.
@@
In the following Methods section notation is first established. Then, supporting (well known) theorems are recalled and the formal constructions of a model, a model function and a model manifold are introduced. The geometry of statistical estimation, in the simple OLS case, is recalled. The problem is then tackled of how the challenges introduced by model nonlinearity in the two settings make the introduction of appropriate tangent spaces necessary. The mapping of the geometry in the random variable space to the geometry of the statistical estimation case space is finally formally defined. A brief Discussion and Conclusions section recaps the results obtained, clarifies the relationship between model manifolds, conditional expectation and information geometry and points to possible further developments.
@@ 2.1.1 Standard notation
Matrices and higher-order arrays are written as (capital) boldface latin letters ($\boldsymbol{H}$ or $H$ when clear from the context). An array may show its dimensions as subscripted integers separated by a cross ($A_{n\times m}$ is the matrix $A$ with $n$ rows and $m$ columns); an array element is identified by its indices (row $r$, column $c$) as $a_{r,c}$ or $a_{rc}$ when no confusion is possible. The $r^{th}$ row and $c^{th}$ column of the matrix $A$ are indicated respectively as $a_{r.}$ and $a_{.c}$   . Notice the difference between $\boldsymbol{x}_{i}$ (the $i^{th}$ vector of a sequence), and $x_{i}$ (the $i^{th}$ scalar of a sequence, usually the $i^{th}$ element of vector $\boldsymbol{x}$).
@@
The real numbers are denoted by $\mathbb{R}$. Positive reals are denoted by $\mathbb{R}^{+}$. Spaces or manifolds are generally indicated by script capital letters like $\mathcal{V}$, $\mathcal{H}$ or $\mathcal{M}$.
Transposition of an array is denoted by superscript capital $T$, like in $J^{\intercal}$. An array, of appropriate dimensions, having all elements equal to 00 (or equal to $1$) is indicated as $\boldsymbol{0}$ (or as $\boldsymbol{1}$ ). The identity matrix of appropriate dimensions is indicated by $\boldsymbol{I}$
The space spanned or generated by a set of vectors is indicated as $\mathcal{S}({\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{m}})$ ; when applied to a matrix, the span function identifies the space generated by the column vectors of the matrix, $\mathcal{S}(A)$
The null space or kernel of a matrix $\boldsymbol{A}$ is denoted by $Ker(\boldsymbol{A})$: $Ker(\boldsymbol{A})=\{\boldsymbol{v}:\boldsymbol{Av}=0\}$.
The orthogonal complement to a linear subspace $V$ is denoted by $V^{\perp}$ ("V-perp"): $V^{\perp}=\{\boldsymbol{x}:\boldsymbol{x}^{\intercal}\cdot\boldsymbol{v}=0\ \ \forall\boldsymbol{v}\in V\}$.
The generalized inverse of a matrix $\boldsymbol{A}$ is indicated as $\boldsymbol{A}^{-}$ and satisfies $\boldsymbol{A}\boldsymbol{A}^{-}\boldsymbol{A}=\boldsymbol{A}$
@@
End of proofs are indicated with the square box ∎.
@@ 2.1.2 Further notation
So far, the notation has been following standard statistical practice. In what follows, however, some necessary further notational conventions are established, which will make it easier to develop the argument in a precise way.
We let $\boldsymbol{x}^{o}=[x^{o}_{i}]_{n\times 1}$ be a vector of observed ("o" stands for observed) realizations of a dependent variable (criterion), possibly corresponding to arbitrarily assigned or anyhow measured values of $m$ independent variables (predictors) $\boldsymbol{U}=[u_{ik}]_{n\times m}=[\boldsymbol{u}_{i.}]_{n\times m}$ , one of which is often time $t$. We also let $\boldsymbol{\theta}=[\theta_{j}]_{q\times 1}$ be a (vector-valued) parameter, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, and we suppose that the (generally nonlinear) functional form of the relationship between criterion and predictors is
@@
In much of our discussion we will not need to indicate the predictors explicitly, since they are part of the experimental design and are fixed. We will thus simply write our predictor function as
@@
We indicate with a "hat" the estimated value of the parameter, $\hat{\boldsymbol{\theta}}$, to which corresponds a "best" prediction or forecast $\hat{\boldsymbol{x}}=\boldsymbol{x(\hat{\boldsymbol{\theta}})}$. We may also suppose that, the postulated relationship being the true one, there exists a "true" value of the parameter, $\boldsymbol{\theta}^{*}$, to which there corresponds the true state of the system $\boldsymbol{x}^{*}$ if it were observed without error.
@@
Using an alternative notation for the Jacobian we may write (in the spirit of Seber [2])
@@
with the obvious meaning being attributed to $\hat{\boldsymbol{\overset{.}{x}}}$, $\hat{\boldsymbol{\overset{..}{x}}}$, $\boldsymbol{\overset{.}{x}}^{*}$ and $\boldsymbol{\overset{..}{x}}^{*}$
Multiplications involving these three-dimensional arrays are defined, for the scope of these notes, simply as the slice-wise equivalent: imagine the array $\boldsymbol{A}_{m\times n\times q}$ to be a pile of $n$ slices, each of which is a $m\times q$ matrix $(\boldsymbol{A}_{.i.})_{m\times q}$ Then pre-multiplication by a matrix $\boldsymbol{B}_{r\times m}$ and post-multiplication by a matrix $\boldsymbol{C}_{q\times s}$ are defined as producing $n$-piles of matrices $(\boldsymbol{BA}_{.i.})_{r\times n\times q}$ or $(\boldsymbol{A}_{.i.}\boldsymbol{C})_{m\times n\times s}$ respectively, with the obvious understanding that dimensions collapsing to $1$ disappear, so that, for example, we treat $\boldsymbol{v}_{1\times m}\boldsymbol{A}_{m\times n\times q}$ simply as the matrix $(\boldsymbol{vA})_{n\times q}$.
@@
  • an orthonormal basis for $\mathbb{R}^{n}$ is a collection of $n$ unit-length, orthogonal basis vectors, say $\boldsymbol{H}=\{\boldsymbol{\eta}_{1},\dots,\boldsymbol{\eta}_{n}\}$ or $\boldsymbol{P}=\{\boldsymbol{\rho}_{1},\dots,\boldsymbol{\rho}_{n}\}$.
  • a vector $\boldsymbol{v}$ is the sum of its components $\{\boldsymbol{v}_{i}\}_{i=1}^{n}$, each of which can be expressed as a coefficient multiplying a basis vector, say $\boldsymbol{v}_{i}=v_{\eta}^{i}\cdot\boldsymbol{\eta}_{i}$, where we emphasize that the coefficient $v_{\eta}^{i}$ refers to the chosen basis $\boldsymbol{H}$
@@
  • we may collect the coefficients into a column array of coefficients, relative to the chosen basis, say $\boldsymbol{v}_{\eta}=\{v^{1}_{\eta},...,v^{n}_{\eta}\}^{T}$    or    $\boldsymbol{v}_{\rho}=\{v^{1}_{\rho},...,v^{n}_{\rho}\}^{T}$
  • we can thus concisely express a vector as its expansion with respect to a given basis:    $\boldsymbol{v}=\boldsymbol{E}\boldsymbol{v}_{\eta}=v_{\eta}^{i}\boldsymbol{\eta}_{i}$    or    $\boldsymbol{v}=\boldsymbol{P}\boldsymbol{v}_{\rho}=v_{\rho}^{i}\boldsymbol{\rho}_{i}$

### v2

@@ Abstract
Mathematical modeling has become pervasive in applications, not only in physics or economy, but also in biomedicine and other “soft” sciences. To the conceptual formulation of a model there often follows its identification by statistical parameter estimation, given available observations. While the nature of the modeling process as well as its relationship with the attending statistical computations could both appear obvious to the practitioner, it may be useful to formalize them in a precise way. Insight into the process of (linear and nonlinear) model parameter estimation can be obtained from the description of the geometry of estimation in case space. The objective then is to also describe the geometry of modeling in the abstract, and to show how the correspondence between the conceptual context and the computational context can be formally represented.
@@ 1 Introduction
Mathematical modeling is central to defining and solving problems in the physical sciences, engineering or economics. Over the past few decades it has found progressively more widespread applications also in biomedicine and other “soft” sciences. In the following we will make use, just by way of example, of naively simple problems in biology: while this choice stems from the background of the author as a biomedical modeler, the mathematical ideas are the same for any of the above application fields.
@@
The initial, modeling phase thus consists in imagining possible relationships among some quantitative features of interest of the reality under study, relationships usually expressed as algebraic or differential equations of various kinds. The initial concern of the present work is to understand exactly what the variables in these equations represent. Suppose for example we wished to estimate weight from height, and we were to write, as a very first attempt to express weight as a function of height, the equation
@@
Notice that at the present stage we are not dealing with any numbers at all, we are at the first stage of idealization of a relationship between a generic concept of height with a generic concept of weight. Our basic question is then: “What, in the above equation, is HH?” The immediate, naïve answer would be “Height!”. Yes, clearly, but whose height? Obviously $H$ is not meant to represent the numerical value of the height of any single individual, to be put in relationship with the weight $W$ of that same individual: changing individuals would then in general change the value of the parameter $k$, which is instead intended to express some general proportion between the height of people and their weight. Similarly, Eq. 1 does not express the relationship between an algebraic variable $W$ and an algebraic variable $H$: we do not imagine people’s weights to follow exactly that functional relationship. Rather, we might consider $H$ as indicating a quantity (height) which could be measured on some individual or on a group of individuals, quite possibly with variability due to heterogeneity among people and/or to measurement error: while we have not yet measured anything at the present stage of the modeling exercise, we could do so in the future. We are thus led to regard $H$ as a random variable.
@@
At this point we consider what $W$ might be. Notice that, if we were hypothetically to reflect on the observable weight of individuals and, following the previous line of reasoning, we were to consider such observable weight of individuals as a random variable, then this random variable (let us denote it provisionally by $W^{o}$) would not be the same as $W$. That $W$ is a random variable follows formally from the fact that it is a function of some numerical value $k$ and of another random variable $H$, so $W$ is indeed a random variable. But $W$ is that random variable obtained from $H$ through some sort of mathematical manipulation, whereas $W^{o}$ is another random variable, representing some characteristic of the people whom we may study in the future, characteristic which may or may not depend on $H$ and, even if it did depend on $H$, it could well in general not depend on it through Eq. 1. In other words, we might consider $W^{o}$ as the objective (hence the o) random variable that we are trying to approximate through some other random variable $W$, which is itself computable from $H$. We are thus led to the definition of a model, which will be formalized in the following, as an operator yielding a random variable (in this case $W$), with which we try to approximate another, objective random variable (in this case $W^{o}$). As will be made more precise in the following, “approximation” has meaning within a space endowed at least with a norm, while other useful concepts, such as projections, will need further structure (inner products), thus leading us to deal with a Hilbert space of random variables as the natural setting for the abstract modeling effort.
@@
If we now consider the second, identification phase of modeling practice, we see that it is commonly performed through the minimization of a cost functional (e.g. by Ordinary Least Squares, OLS, or more generally by Maximum Likelihood, ML). Identification of course depends on having sampled some individuals from the population of interest and having observed on them both the intended criterion or target variable and the intended predictor variable(s) (such as, in the previous example, both $W^{o}$ and $H$, from which last realizations of $W$ may be calculated). This procedure, which is inherently computational, is executed within an $n$-dimensional Euclidean space, referred to here as the case space.
While both phases of the model construction process, modeling and identification, have been conducted innumerable times over centuries and are being more and more intensively conducted by applied mathematicians all over the world, to the best of the author’s knowledge the relationship between abstract model building and concrete statistical estimation computations has so far been only intuited as their "obvious" correspondence. The aim of the present work is to make formally precise the relationship between the ideal construction of a mathematical model in the space of random variables and the statistical estimation procedure in $n$-dimensional case space.
@@
It should be noted that the consideration of the Hilbert space $\mathcal{H}$ of (finite-variance) random variables is not new, see for example Small [Small1994]. However, subsequent developments have centered mainly on the characterization of conditional expectation as the projection onto a linear subspace of $\mathcal{H}$ [Bobrowski2013], generated by an appropriate sub-$\sigma$-algebra of the original probability space. The present work, instead, focuses on the modeling process itself, giving rise to a generally nonlinear model manifold, and on the relationship of this construction with the (geometrical interpretation of) commonly performed statistical parameter estimation procedures.
The following discussion links more closely with the subject matter of information geometry [Amari2009, Amari2010, BarndorffNielsen1986], which deals with manifolds of probability distributions. While there exists a correspondence between (equivalence classes of) random variables and their induced probability distributions, and while differential geometric methods in information geometry help establish properties (e.g. of convergence) of families of probability distributions, the focus here is on the direct expression of the model structure itself as a manifold in a space of random variables. In other words, while there exists a correspondence, the spaces dealt with in information geometry and here are different (one a space of distributions, the other a space of variables).
@@
In the following, notation is first established. Then, supporting (well known) theorems are recalled and the formal constructions of a model, of a model function and of a model manifold are introduced. The geometry of statistical estimation, in the simple OLS case, is recalled. The problem is then tackled of how the challenges introduced by nonlinearity of the model in the two settings make the introduction of appropriate tangent spaces desirable. The mapping of the geometry in the random variable space to the geometry of the statistical estimation case space is finally formally defined. A brief final section recaps the results obtained, clarifies the relationship between model manifolds, conditional expectation and information geometry and points to possible further developments.
@@ 2.1.1 Standard notation
Matrices and higher-order arrays are written as (capital) boldface latin letters ($\boldsymbol{H}$ or $H$ when clear from the context). An array may show its dimensions as subscripted integers separated by a cross ($A_{n\times m}$ is the matrix $A$ with $n$ rows and $m$ columns); an array element is identified by its indices (row $r$, column $c$) as $a_{r,c}$ or $a_{rc}$ when no confusion is possible. The $r^{th}$ row and $c^{th}$ column of the matrix $A$ are indicated respectively as $a_{r.}$ and $a_{.c}$   . Notice the difference between $\boldsymbol{x}_{i}$ (the $i^{th}$ vector of a sequence), and $x_{i}$ (the $i^{th}$ scalar of a sequence, such as the $i^{th}$ element of vector $\boldsymbol{x}$).
@@
The real numbers are denoted by $\mathbb{R}$. Non-negative reals are denoted by $\mathbb{R}^{+}$. Spaces or manifolds are generally indicated by script capital letters like $\mathcal{V}$, $\mathcal{H}$, $\mathcal{M}$ or $\mathcal{Y}$.
Transposition of an array is denoted by superscript capital $T$, like in $J^{\intercal}$. An array, of appropriate dimensions, having all elements equal to 0 (or equal to $1$) is indicated as $\boldsymbol{0}$ (or as $\boldsymbol{1}$ ). The identity matrix of appropriate dimensions is indicated by $\boldsymbol{I}$ .
The space spanned or generated by a set of vectors is indicated as
$\mathcal{S}({\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{m}})$; when applied to a matrix $A$ the span function $\mathcal{S}(A)$ indicates the space generated by the column vectors of the matrix.
The null space or kernel of a matrix $\boldsymbol{A}$ is denoted by $Ker(\boldsymbol{A})$,   $Ker(\boldsymbol{A})=\{\boldsymbol{v}:\boldsymbol{Av}=0\}$.
The orthogonal complement to a linear subspace $\mathcal{V}$ is denoted by $\mathcal{V}^{\perp}$ ($\mathcal{V}$-perp"),   $\mathcal{V}^{\perp}=\{\boldsymbol{x}:\boldsymbol{x}^{\intercal}\cdot\boldsymbol{v}=0\ \ \forall\boldsymbol{v}\in\mathcal{V}\}$.
The generalized inverse of a matrix $\boldsymbol{A}$ is indicated as $\boldsymbol{A}^{-}$ and satisfies $\boldsymbol{A}\boldsymbol{A}^{-}\boldsymbol{A}=\boldsymbol{A}$ .
@@
End of proofs are indicated with the square box .
@@ 2.1.2 Further notation
So far, the notation reflects standard statistical practice. In what follows, however, some necessary further notational conventions are established, which will make it easier to develop the argument in a precise way.
We let $\boldsymbol{x}^{o}=[x^{o}_{i}]_{n\times 1}$ be a vector of observed ("o" stands for observed) realizations of a dependent variable (criterion), possibly corresponding to arbitrarily assigned or measured (assumedly without error) values of $m$ independent variables (predictors) $\boldsymbol{U}=[u_{ik}]_{n\times m}=[\boldsymbol{u}_{i.}]_{n\times m}$ , one of which is often time $t$. We also let $\boldsymbol{\theta}=[\theta_{j}]_{q\times 1}$ be a (vector-valued) parameter, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{q}$, and we suppose that the (generally nonlinear) functional form of the relationship between criterion and predictors is
@@
In much of our discussion we will not need to indicate the predictors explicitly, since they are part of the experimental design and are fixed. With a slight abuse of notation we will thus simply write our predictor function as
@@
We indicate with a "hat" the estimated value of the parameter, $\hat{\boldsymbol{\theta}}$, to which corresponds a "best" prediction or forecast $\hat{\boldsymbol{x}}=\boldsymbol{x(\hat{\boldsymbol{\theta}})}$. We also suppose that, the postulated relationship being the true one, there exists a "true" value of the parameter, $\boldsymbol{\theta}^{*}$, to which corresponds the true state of the system $\boldsymbol{x}^{*}=\boldsymbol{x}(\boldsymbol{\theta}^{*})$.
@@
Using an alternative notation for the Jacobian we may write (in the spirit of Seber [Seber2003])
@@
with the obvious meaning being attributed to $\hat{\boldsymbol{\overset{.}{x}}}$, $\hat{\boldsymbol{\overset{..}{x}}}$, $\boldsymbol{\overset{.}{x}}^{*}$ and $\boldsymbol{\overset{..}{x}}^{*} .$
Multiplications involving these three-dimensional arrays are defined, for the scope of the present work, simply as the slice-wise equivalent: the array $\boldsymbol{A}_{m\times n\times q}$ is treated as a pile of $n$ slices, such that the $i$-th slice is an $m\times q$ matrix $(\boldsymbol{A}_{.i.})_{m\times q}$ . Then pre-multiplication by a matrix $\boldsymbol{B}_{r\times m}$ and post-multiplication by a matrix $\boldsymbol{C}_{q\times s}$ are defined as producing $n$-piles of matrices $(\boldsymbol{BA}_{.i.})_{r\times n\times q}$ or $(\boldsymbol{A}_{.i.}\boldsymbol{C})_{m\times n\times s}$ respectively, with the understanding that dimensions collapsing to $1$ disappear, so that, for example, we treat $\boldsymbol{v}_{1\times m}\boldsymbol{A}_{m\times n\times q}$ simply as the matrix $(\boldsymbol{vA})_{n\times q}$.
@@
  • A Hamel basis of a vector space $\mathcal{V}$ over a field $\mathbb{K}$ is a set $\mathcal{B}\subset\mathcal{V}$ such that every element $v\in\mathcal{V}$ can be written uniquely as a finite linear combination of elements of $\mathcal{B}$. This means that $\mathcal{B}$ is both linearly independent and spans $\mathcal{V}$ in the sense that for every $v\in\mathcal{V}$, there exist finitely many $b_{i}\in\mathcal{B}$ and scalars $\alpha_{i}\in\mathbb{K}$ such that $v=\sum_{i=1}^{n_{v}}\alpha_{i}b_{i}$.
  • In finite-dimensional spaces (such as $\mathbb{R}^{d}$) the standard basis with $d$ elements is a Hamel basis. However, in infinite-dimensional spaces, Hamel bases are typically uncountable and non-constructive, relying on the Axiom of Choice for their existence.
  • A Schauder basis allows infinite (typically countable) linear combinations. It must be defined in the context of topological vector spaces, where convergence of the series is required.
  • An orthonormal basis for an $n$-dimensional Hilbert space (such as $\mathbb{R}^{n}$) is a collection of $n$ unit-length, orthogonal basis vectors, say $\boldsymbol{H}=\{\boldsymbol{\eta}_{1},\dots,\boldsymbol{\eta}_{n}\}$ or $\boldsymbol{P}=\{\boldsymbol{\rho}_{1},\dots,\boldsymbol{\rho}_{n}\}$.
  • a vector $\boldsymbol{v}$ is the sum of its components $\{\boldsymbol{v}_{i}\}_{i=1}^{n}$, each of which can be expressed as a coefficient multiplying a basis vector, say $\boldsymbol{v}_{i}=\text{v}_{\eta}^{i}\cdot\boldsymbol{\eta}_{i}$, where we emphasize that the coefficient $\text{v}_{\eta}^{i}$ refers to the chosen basis $\boldsymbol{H}$
@@
  • we may collect the coefficients into a column array of coefficients, relative to the chosen basis, say ${\text{{v}}}_{\eta}=\{\text{v}^{1}_{\eta},...,\text{v}^{n}_{\eta}\}^{\intercal}$    or    ${\text{{v}}}_{\rho}=\{\text{v}^{1}_{\rho},...,\text{v}^{n}_{\rho}\}^{\intercal}$
  • we can thus concisely express a vector as its expansion with respect to a given basis:    $\boldsymbol{v}=\boldsymbol{E}{\text{{v}}}_{\eta}=\text{v}_{\eta}^{i}\boldsymbol{\eta}_{i}$    or    $\boldsymbol{v}=\boldsymbol{P}{\text{{v}}}_{\rho}=\text{v}_{\rho}^{i}\boldsymbol{\rho}_{i}$